OpenStudy (anonymous):

Evaluate 4 over 7 - 7 over 15 + 1 over 3

OpenStudy (anonymous):

confusing or not, that is what you have to do

OpenStudy (anonymous):

These are fractions

OpenStudy (anonymous):

yeah i know lets take them one at a time

OpenStudy (anonymous):

\[\frac{4}{7}-\frac{7}{15}\] is the first one

OpenStudy (anonymous):

Ops the first one is 4/5

OpenStudy (anonymous):

4/5 - 7/15 + 1/3 ?

OpenStudy (anonymous):

Yaa

OpenStudy (anonymous):

then it is easier

OpenStudy (anonymous):

Be right back sorry

OpenStudy (anonymous):

It's so simple.. 4/5 = 12/15 ? Understood. 1/3 = 5/15 ?understood. The answer is 12/15+5/15 - 7/15 = 10/15 = 2/3

OpenStudy (anonymous):

I don't get it @vipvipindev

OpenStudy (anonymous):

Understood 4/5 = 12/15 ?

OpenStudy (anonymous):

Noo

OpenStudy (anonymous):

How.... it is bc you simply that to get 4/5

OpenStudy (whpalmer4):

4/5 *3/3=12/15

OpenStudy (anonymous):

For addition we require a need to make the denominator same. So by multiplying 5 with 3 we get it as 15. When we multiply to denominator we also do it for numerator. Understood?

OpenStudy (anonymous):

Ohokkk yes

OpenStudy (anonymous):

Sameway did here : 1/3* 5/5 ?

OpenStudy (anonymous):

\[\frac{ 4 }{ 7 }-\frac{7}{15}+\frac{1}{3}\] Find the common factor of the denominators: 7: 7, 14, 21, 28, 35, 42, 49, 56, 63, 70, 77, 84, 91, 98, 105 15: 15, 30, 45, 60, 75, 90, 105 3: 3, 6, 9, 12, 15, ... , 105 (because 3 x 35 = 105) Multiply the fractions to make the numerator the same as the common factor: \[\frac{ 4 }{ 7} * 15 = \frac{ 60 }{ 105 }\] \[\frac{ 7 }{ 15} * 7 = \frac{ 49 }{ 105 }\] \[\frac{ 1 }{ 3} * 35 = \frac{ 35 }{ 105 }\] Lastly, add/subtract the fractions with the same denominator: \[\frac{ 60 }{105}-\frac{49}{105}+\frac{35}{105} =\frac{46}{105}\]

OpenStudy (anonymous):

So u have to get he the nominator to all be the same

OpenStudy (anonymous):

Denominator **

OpenStudy (whpalmer4):

Yes that is correct

OpenStudy (whpalmer4):

Argilly showed the whole process for the problem as originally posted

OpenStudy (anonymous):

But it says the answer is 2/3

OpenStudy (whpalmer4):

Key is to find the common denominator. One way is to just multiply all of the denominators together. This works, but gives a bigger number than neede if one denominators is a multiple of another.

OpenStudy (anonymous):

I think it's better to give 15 as a common denominator so that he/she can understand!@whpalmer4

OpenStudy (anonymous):

Ohokkkkkk I get it now ... but vipvipindev if e denominator is 15 how did u get 3

OpenStudy (whpalmer4):

\[\frac{4}{5}-\frac{7}{15}+\frac{1}{3}=\frac{4*3}{5*3}-\frac{7}{15}+\frac{1*5}{3*5}=\frac{12-7+5}{15}=\frac{10}{15}=\frac{2*5}{3*5}=\frac{2}{3}\]

OpenStudy (whpalmer4):

@vip hadn't finished yet, doing latex on iPad keyboard while walking is very slow...

OpenStudy (whpalmer4):

You can also find least common multiple to use as common denominator by making lists of multiples to see which comes up first: 3 6 9 12 15 5 10 15 15 15 is first number to appear in all 3 lists so it is least common multiple of 3,5,15

OpenStudy (anonymous):

Ohhh okk I see now...thanks you

OpenStudy (whpalmer4):

Can also do it by factoring denominators and combining 3 factors to 3 5 factors to 5 15 factors to 3*5 We multiply each unique factor found, using as many copies found in the number that had the most. Here that would be 3*5=15

OpenStudy (anonymous):

I get it now wow that was easy now that I noe it

OpenStudy (whpalmer4):

If doing problem with different denominators of 2 8 12 (for example only) 2: 2 8: 2*2*2 12:2*2*3 Here we multiply 2*2*2*3=24 because 2 and 3 were factors found and we had 3 2s in factors for 8 2 4 6 8 10 12 14 16 18 20 22 24 8 16 24 12 24 Again 24 is first number in each list so it is least common multiple of 2,8,12 This example shows how just multiplying denominators together may give unnecessarily large denominator: 2*8*12=192 which is 8x as big as needed!